• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J.P. Chen

    Log entry time 10:25:23 on October 28,1998

    Entry number 3932

    keyword=Target polarization measurements

    NMR measurements for transverse pumping:
    8:45-9:00 turned off transverse lasers, moved into pick-position, and rotated field from 270 to 0 degree. Did NMR with lasers off:
    Peak=12.06 mV, pol=31.4% (use 2.6%/mv calibration).
    9:05 turned on 3 longitudinal lasers, did NMR with lasers on:
    peak=12.33 mV, pol=33.2%.
    After measurements, turned off longitudinal lasers, moved back to pol he3 position, rotated back to 270 degree, put transverse lasers on one by one.